jhonyy9 (jhonyy9):

correct solved ? cosec x lim ln((-------- )/cot x)x = x->e sec x 1/sin x lim ln((--------- )/cot x)x = x->e 1/cos x cos x lim ln(------- )/cot x)x = x->e sin x lim ln(cot x /cot x)x = x->e lim ln x = x->e lim ln e = 1 x->e

jhonyy9 (jhonyy9):

@Kainui

jhonyy9 (jhonyy9):

@TheSmartOne @zepdrix

OpenStudy (agent0smith):

Looks good.

jhonyy9 (jhonyy9):

ty

OpenStudy (sshayer):

\[\lim_{x \rightarrow e}\ln \frac{ \cot x }{ \cot x }*x=\lim_{x \rightarrow e}\ln 1*x\] is it 1*x or (ln1)x ?

OpenStudy (sshayer):

if it is 1*x then limit=1 otherwise ln1=0 0*x=0 then limit is 0

jhonyy9 (jhonyy9):

so there are parentheses ln(----- inclusive *x)

jhonyy9 (jhonyy9):

ok. @sshayer thank you

OpenStudy (sshayer):

yw